1. A 0.3126 g sample of oxalic acid, H2C2O4, requires 26.21 mL of a particular concentration of NaOH(aq) to complete the following reaction. What is the molarity of the NaOH(aq)? Write the full computation of your answer.


H2C2O4(s) + NaOH(aq) → Na2CsO4(aq) + H2O(l)


1
Expert's answer
2021-04-28T02:46:03-0400

H2C2O4(s) + 2NaOH(aq) → Na2C2O4(aq) + 2H2O(l)

From the equation, 1 mole of oxalic acid requires 2 moles of NaOH.

n = m / M

M (H2C2O4) = 90.03 g/mol

n (H2C2O4) = 0.3126 / 90.23 = 0.0035 mol


n (NaOH) = 2 x 0.0035 = 0.007 mol

CM (NaOH) = n / V = 0.007 / 0.02621 = 0.26 M